One number is 1 more than 2 times another. Their product is 36. What are those two numbers

Found 2 solutions by checkley79, richwmiller:
Answer by checkley79(3341)

X=2Y+1
X*Y=36
(2Y+1)Y=36
2Y*2+Y=36
2Y^2+Y-36=0
(2Y-8)(Y+4.5)=0
2Y-8=0
2Y=8
Y=8/2
Y=4 ANS.
X=2*4+1=8+1=9 ANS.
PROOF:
8*4=36
36=36

Answer by richwmiller(17219)

the check of the other tutor is incorrect.
9*4=36